The problem is -(1/5)d-1=12
Answer by checkley77(12844)

-(1/5)d-1=12
-d/5-1=12
-d/5=12+1
-d/5=13
-d=65
d=-65
proof:
-1/5*-65-1=12
65/5-1=12
13=12+1
13=13